(a) Your company’s current ratio is 0.5x, while your competitor’s current ratio is 1.5x. Both
firms want to "window dress" the coming end-of-year financial statements. As part of
the window dressing strategy, each firm will double its current liabilities by adding
short-term debt and placing the funds obtained in the cash account. Describe the actual
results of these transactions?

(b) What would be the effect if a company increases its debt ratio, but leaves its operating
income (EBIT) and total assets unchanged?

(c) A fire has destroyed a large percentage of the financial records of the Sunlight Inc. You
have the task of piecing together information in order to release a financial report. You
have found the return on equity to be 18 percent. If sales were RM4 million, the debt
ratio was 0.40, and total liabilities were RM2 million, what was the return on assets
(ROA)?
1
Expert's answer
2016-03-18T15:48:04-0400
(a) The actual results of these transactions will be better for your competitor, because now your current debts are 3 times higher, then your current liabilities, which creates some risks for your company.
(b) If a company increases its debt ratio, but leaves its operating income (EBIT) and total assets unchanged, the share of debt will increase without any improvements in other indicators, so the risks will be increased.
(c) ROE = 18 percent, sales = RM4 million, the debt ratio = 0.40, and total liabilities = RM2 million, then total assets = total liabilities/debt ratio = 2/0.4 = RM 5 million. Equity = Total assets - Total liabilities = 5 - 2 = RM3 million. ROE = Net income/Equity = 0.18, so Net income = 0.18*3 = RM0.54 million.
The return on assets ROA = Net income/Total assets = 0.54/5 = 0.108